# Light-ResKAN: A Parameter-Sharing Lightweight KAN with Gram Polynomials for Efficient SAR Image Recognition

### 02 Apr 2026 v1

@@ I Introduction
Synthetic Aperture Radar (SAR) is a radar system that produces high-resolution imagery for Earth observation by transmitting electromagnetic pulses and processing the returned signals[Wiley1985]. Capable of all-weather and all-day operation, SAR actively supports extensive applications in environmental monitoring and defense, spanning hydrology, topography, resource management, and disaster response [ref1, xu2024cloudseg, jiang2025high]. SAR image recognition is the technology for automated detection and identification of various targets and patterns in SAR imagery through computational methods. In recent years, the rapid advancements in deep learning technology have offered technical avenues for SAR target recognition [Fein-Ashley_2023, li2024predicting, LiSARATRX25] and various deep neural networks have significantly propelled research progress in this domain through their proficient feature extraction capabilities [datcu2023explainable, hua2015deep, 726791, 10533864, hinton2006reducing, 11146877, 11072100, williams2013gradient, goodfellow2014generative, vaswani2017attention, liu2024s, Liu2025Causal]. These methods learn deep feature representations from images, enabling high-accuracy classification and detection tasks [Yan_2024_11, zhou2024madinet, 10977792, zhou2024diffdet4sar].
@@
However, their substantial model complexity and high computational resource demands present significant limitations for deployment in resource-constrained environments, such as unmanned aerial vehicles (UAVs) and satellites [Housseini_2017, zhou2025fiftyyearssaratr]. Concurrently, SAR images captured by UAVs or satellite platforms often exhibit larger spatial extents and pixel dimensions compared to conventional RGB imagery. This characteristic, stemming from the need for wide geographical coverage, leads to a rapid escalation in computational workload and resource consumption. Consequently, the development of lightweight models for SAR image recognition has become a paramount research focus.
@@
Conventional model compression techniques—including network pruning, knowledge distillation, quantization and low-rank approximation—have been widely explored [Li_2023, Zhou_2025_04, su2023lightweight, yang2024dynamic, Choudhary_2020_02, Li_2023_03]. A dominant trend involves drawing inspiration from lightweight architectures such as MobileNet [howard2019searching] and ShuffleNet [ma2018shufflenet], which heavily employ depthwise separable convolutions to enhance parameter and computational efficiency. Another line of work adapts mechanisms from vision transformers, such as the hierarchical window attention in Swin Transformer [liu2021swin], to significantly reduce computational overhead. Therefore, mainstream research has primarily focused on refining key components within established CNN and Transformer architectures. Nevertheless, although they achieved significant results in reducing model size and computational burden, they often come at the cost of sacrificing recognition accuracy due to the loss of high-frequency information caused by model compression, limiting overall performance and robustness [18, 21]. This inherent trade-off presents a persistent challenge in deploying efficient yet highly performant deep learning models in resource-constrained environments.
@@
The emerging Kolmogorov-Arnold Network (KAN) [7] enhances model expressiveness and interpretability through learnable activation functions, fundamentally adhering to the mathematical principles of Kolmogorov’s superposition theorem. Theoretically, this architecture requires only 2n+1 activation functions to exactly approximate arbitrary continuous functions. Compared to conventional deep learning frameworks, KAN demonstrates a substantial parameter efficiency advantage. While maintaining nonlinear approximation capabilities, it is more parameter-efficient than traditional MLPs and circumvents the necessity for stacked convolutions in CNNs and the quadratic growth in computational complexity inherent in Transformer architectures. This architectural innovation has been widely implemented in remote sensing image recognition tasks, demonstrating practical efficacy across diverse application scenarios [240, 25, 26, KAN_ECCV2024]. The ability of KANs to model complex, non-linear relationships efficiently makes them a strong candidate for tackling the intricate patterns present in SAR data for resource-constrained environments [Granata2024].
@@ II-A3 Transformer-Based Lightweight Models
While most work relies on CNNs, Zhao et al. [zhao24] explored Vision Transformers (ViTs) for SAR processing, proposing a Lightweight Visual Transformer (LViT) that encodes images via a ViT backbone and processes features through an MLP for efficient inference.
Although the above methods have made remarkable progress in improving the efficiency and accuracy of target recognition in SAR images, they each have some limitations. Although the depth separable convolution used in [18] and [20], [21] and [23] effectively reduce the computational cost, it is not enough to deal with the extraction of high-precision features in a complex electromagnetic environment, and it is easy to lose the unique high-frequency scattering features of SAR images. The proposed R-RetinaNet by [19] mitigates the conflict between learning objectives by decoupling the feature optimization process for different tasks. However, this approach may increase model complexity and may not be friendly to resource-constrained edge devices. The method of [22] relies on the fine-tuning of the MobilenetV2 architecture of the self-attention mechanism as a binary classifier, which improves the classification accuracy to some extent, but may lack generality and flexibility for multiclass SAR image recognition problems. [zhao24] uses the ViT encoder combined with MLP to realize the lightweight visual transformer (LViT). Although it has advantages in handling long-distance dependencies, it still faces challenges in real-time and computational efficiency.
These specific shortcomings reflect broader, more fundamental limitations of conventional CNNs and Transformers in modeling the complex nonlinear mappings required for SAR imagery. While CNNs excel at capturing local spatial features, their extensive stacking of convolutional layers often leads to prohibitive computational complexity and increased memory burden. Transformers, though effective in global dependency modeling, suffer from the quadratic complexity of their self-attention mechanism, which becomes a significant bottleneck for high-resolution SAR data. Furthermore, both architectures typically require extensive hyperparameter tuning, custom architectural adjustments, and protracted retraining cycles [Xu_2025_07, Rakesh_2025], which can undermine their generalizability and robustness against SAR-specific challenges like speckle noise and significant resolution variations [Lee2023ResourceEfficient]. This analysis underscores a pressing need for novel, inherently lightweight architectural designs that possess stronger local feature extraction capabilities, superior parameter efficiency, and enhanced nonlinear expressive power beyond the conventional paradigms.
@@ II-B2 Application of KAN in Remote Sensing Field
Minjong Cheon [240] integrated KAN with various pre-trained CNN models to obtain KCN for remote sensing (RS) scene classification tasks. [25] proposed SpectralKAN, which can achieve high hyperspectral image change detection (HSIs-CD) accuracy, while requiring less parameters, FLOPs, GPU memory, training and testing time, and improving the efficiency of the HSIs-CD. [26] proposed a Wavelet-based Kolmogorov-Arnold Network Architecture (WaveKAN) that incorporates Wavelet functions into learnable activation functions for hyperspectral image classification. [KAN_ECCV2024] integrated the KAN layer into the U-Net architecture (U-KAN), segmented farmland using Sentinel-2 and Sentinel-1 satellite images, and analyzed the performance and interpretability of these networks. Table II is a summary of the above methods.
@@ III-A The Design Motivation of Light-ResKAN
SAR images exhibit distinct characteristics compared to optical imagery, owing to its unique imaging mechanism and data properties [Guo_2024]. Specifically, SAR images are severely afflicted by coherent speckle noise during the imaging process. Furthermore, due to the influence of side-looking imaging geometric distortions, their spatial resolution is typically lower, commonly ranging between 0.1 and 10 meters [Luo_2023], [Zhang_2023_01]. This resolution is notably insufficient when contrasted with optical imagery acquired from the same platform.
@@ III-C1 Shared Activation Function Convolutions
As an orthogonal polynomial sequence, Gram Polynomials exhibit theoretical properties that align exceptionally well with the unique challenges of SAR imagery, positioning them as a highly promising activation function for SAR image recognition. Their principal advantage lies in their ability to adaptively integrate the processes of feature extraction and speckle noise suppression through adjustable polynomial degrees—effectively mitigating coherent speckle in homogeneous regions while enhancing critical features at edges and strong scatterers. Meanwhile, they adeptly handle the high dynamic range inherent in SAR data with their oscillatory characteristics. The orthogonality of their basis also contributes to numerical stability and generalization capability. More importantly, as demonstrated by the pioneering work of Liu et al. [Liu2016MRELBP] in texture classification, polynomial like features can effectively capture the texture structure information of images. In addition, research on successful texture classification from random features [liuRandomFeatures] also indicates that methods based on function approximation can demonstrate significant strong potential in texture representation tasks. Consequently, from a theoretical perspective, Gram Polynomials offer a more powerful and intelligent nonlinear transformation compared to fixed functions like ReLU, providing a novel solution for SAR image recognition that combines mathematical rigor with enhanced practical adaptability.
@@ IV-B Experimental Setup
AlexNet [alexnet] 2012 57.04 228.18 97.44 89.41 95.76 0.19 3.71 0.19 9.54 30.32 9.53
VGG16 [vgg] 2015 134.30 537.21 96.59 90.73 96.17 4.07 80.37 4.07 11.42 43.72 11.40
ResNet18 [resnet] 2016 11.18 44.73 96.71 83.54 95.34 0.49 9.51 0.49 10.31 32.41 10.29
ResNet50 [resnet] 2016 23.53 94.11 97.13 86.71 96.69 1.08 21.35 1.08 10.92 32.93 10.91
ConvNeXt [convnext] 2022 87.58 350.23 96.04 77.86 95.83 4.07 80.53 4.07 17.23 53.25 17.19
MACN [macn] 2022 0.05 0.16 82.46 89.15 79.30 0.63 12.40 0.63 97.46 296.75 97.46
Cross-HL [cross-hl] 2024 0.52 2.06 98.19 78.06 97.03 0.01 0.22 0.01 11.57 35.64 11.56
ASANet [asanet] 2024 27.95 111.28 91.65 78.84 96.81 2.37 46.82 2.37 18.42 47.23 18.41
@@ IV-C Experimental Results
In this experiment, the performance of Light-ResKAN is comprehensively compared with a variety of representative models on the FUSAR-Ship, MSTAR, and SAR-ACD datasets. These baselines include: (1) Classic CNN architectures such as AlexNet [alexnet], VGG16 [vgg], and the ResNet series (ResNet18, ResNet50) [resnet], which established the foundation for deep feature extraction. (2) ConvNeXt [convnext], a modernized pure CNN that integrates Transformer-like design principles. (3) Lightweight models specifically designed for efficiency, including MACN [macn] which uses depthwise separable convolutions and Cross-HL [cross-hl] featuring cross-layer feature fusion. (4) ASANet [asanet], a multi-scale attention network specialized for SAR and RGB image recognition. (5) KAN-based competitors including SpectralKAN [25], WaveKAN [26], and the original ResKAGNet [8] which serves as the direct baseline for our proposed lightweight framework. Table VI shows the test accuracy, parameter quantity, model size, FLOPs per batch, and test time per batch of different models on three SAR dataset. The best results in each column (highest accuracy, lowest parameters, model size, FLOPs, and test time) are highlighted in bold.
@@
Test time. AlexNet yields the shortest test time due to its simple and highly optimized operator pipeline. Despite Light-ResKAN achieving superior accuracy with significantly reduced parameters and FLOPs, it currently exhibits higher inference latency compared to ResKAGNet and other lightweight baselines. This discrepancy highlights the disconnect between theoretical complexity metrics and actual runtime speed, a phenomenon comprehensively analyzed in ShuffleNet-V2 [ma2018shufflenet]. The observed latency primarily stems from two critical factors: the inherent computational characteristics of KANs and the specific engineering constraints necessitated by our implementation.
@@ V Discussions
AlexNet [alexnet] 99.38 15.01 86.31
VGG16 [vgg] 99.34 320.83 131.50
ResNet18 [resnet] 99.05 38.01 90.47
ResNet50 [resnet] 99.17 86.58 101.47
ConvNeXt [convnext] 98.76 321.47 166.58
MACN [macn] 92.53 49.48 956.23
Cross-HL [cross-hl] 99.43 0.84 89.92
ASANet [asanet] 95.36 186.81 145.93
@@ V-C1 Noise Characteristics of SAR Images
SAR systems utilize coherent radar technology, transmitting coherent pulses and performing coherent superposition of the echo signals upon reception [chen2024very]. When the radar beam illuminates the ground, a resolution cell contains multiple scatterers. The echoes from these scattering points are superimposed vectorially at the receiver. Due to the independent and random distribution of amplitudes and phases of the individual scatterers, the superimposed total echo intensity exhibits random fluctuations on top of the scattering coefficient, forming alternating bright and dark speckle noise. This superposition process directly leads to the noise being correlated with the local intensity of the signal, manifesting as multiplicative noise [twinkle2024edge]. The specific formula is as follows:
@@
Where $I$ represents the observed value, $R$ represents the ideal radar scattering cross-section unaffected by noise, reflecting the backscattering characteristics of the ground object, and $F$ represents the coherent speckle noise component. To suppress noise, SAR often employs multi-look processing, which involves averaging multiple independent observations of the same area [gallet2024renyi]. According to the Central Limit Theorem, the intensity noise after multi-look processing obeys a Gamma distribution, and the probability density function of its noise variable $F$ is:
@@
Where $\alpha$ is the shape parameter that determines the distribution shape, and $\beta$ is the scale parameter that controls the distribution’s scale. Additionally, studies have shown that the histogram of SAR image intensity data fits the Gamma distribution better than the Gaussian distribution, especially in low-resolution or complex scenes, where the gamma model can more accurately describe the clutter statistical properties [wang2024novel]. Based on this, we add multiplicative noise following a Gamma distribution to three SAR datasets.

### 03 Apr 2026 v2

@@ I Introduction
Synthetic Aperture Radar (SAR) is a radar system that produces high-resolution imagery for Earth observation by transmitting electromagnetic pulses and processing the returned signals[53]. Capable of all-weather and all-day operation, SAR actively supports extensive applications in environmental monitoring and defense, spanning hydrology, topography, resource management, and disaster response [47, 57, 18]. SAR image recognition is the technology for automated detection and identification of various targets and patterns in SAR imagery through computational methods. In recent years, the rapid advancements in deep learning technology have offered technical avenues for SAR target recognition [8, 24, 23] and various deep neural networks have significantly propelled research progress in this domain through their proficient feature extraction capabilities [5, 17, 20, 56, 14, 65, 45, 54, 10, 50, 32, 28]. These methods learn deep feature representations from images, enabling high-accuracy classification and detection tasks [60, 70, 67, 71].
@@
However, their substantial model complexity and high computational resource demands present significant limitations for deployment in resource-constrained environments, such as unmanned aerial vehicles (UAVs) and satellites [7, 69]. Concurrently, SAR images captured by UAVs or satellite platforms often exhibit larger spatial extents and pixel dimensions compared to conventional RGB imagery. This characteristic, stemming from the need for wide geographical coverage, leads to a rapid escalation in computational workload and resource consumption. Consequently, the development of lightweight models for SAR image recognition has become a paramount research focus.
@@
Conventional model compression techniques—including network pruning, knowledge distillation, quantization and low-rank approximation—have been widely explored [22, 72, 46, 63, 4, 25]. A dominant trend involves drawing inspiration from lightweight architectures such as MobileNet [16] and ShuffleNet [37], which heavily employ depthwise separable convolutions to enhance parameter and computational efficiency. Another line of work adapts mechanisms from vision transformers, such as the hierarchical window attention in Swin Transformer [33], to significantly reduce computational overhead. Therefore, mainstream research has primarily focused on refining key components within established CNN and Transformer architectures. Nevertheless, although they achieved significant results in reducing model size and computational burden, they often come at the cost of sacrificing recognition accuracy due to the loss of high-frequency information caused by model compression, limiting overall performance and robustness [43, 29]. This inherent trade-off presents a persistent challenge in deploying efficient yet highly performant deep learning models in resource-constrained environments.
@@
The emerging Kolmogorov-Arnold Network (KAN) [35] enhances model expressiveness and interpretability through learnable activation functions, fundamentally adhering to the mathematical principles of Kolmogorov’s superposition theorem. Theoretically, this architecture requires only 2n+1 activation functions to exactly approximate arbitrary continuous functions. Compared to conventional deep learning frameworks, KAN demonstrates a substantial parameter efficiency advantage. While maintaining nonlinear approximation capabilities, it is more parameter-efficient than traditional MLPs and circumvents the necessity for stacked convolutions in CNNs and the quadratic growth in computational complexity inherent in Transformer architectures. This architectural innovation has been widely implemented in remote sensing image recognition tasks, demonstrating practical efficacy across diverse application scenarios [3, 52, 42, 39]. The ability of KANs to model complex, non-linear relationships efficiently makes them a strong candidate for tackling the intricate patterns present in SAR data for resource-constrained environments [11].
@@ II-A3 Transformer-Based Lightweight Models
While most work relies on CNNs, Zhao et al. [68] explored Vision Transformers (ViTs) for SAR processing, proposing a Lightweight Visual Transformer (LViT) that encodes images via a ViT backbone and processes features through an MLP for efficient inference.
Although the above methods have made remarkable progress in improving the efficiency and accuracy of target recognition in SAR images, they each have some limitations. Although the depth separable convolution used in [43] and [30], [29] and [58] effectively reduce the computational cost, it is not enough to deal with the extraction of high-precision features in a complex electromagnetic environment, and it is easy to lose the unique high-frequency scattering features of SAR images. The proposed R-RetinaNet by [62] mitigates the conflict between learning objectives by decoupling the feature optimization process for different tasks. However, this approach may increase model complexity and may not be friendly to resource-constrained edge devices. The method of [1] relies on the fine-tuning of the MobilenetV2 architecture of the self-attention mechanism as a binary classifier, which improves the classification accuracy to some extent, but may lack generality and flexibility for multiclass SAR image recognition problems. [68] uses the ViT encoder combined with MLP to realize the lightweight visual transformer (LViT). Although it has advantages in handling long-distance dependencies, it still faces challenges in real-time and computational efficiency.
These specific shortcomings reflect broader, more fundamental limitations of conventional CNNs and Transformers in modeling the complex nonlinear mappings required for SAR imagery. While CNNs excel at capturing local spatial features, their extensive stacking of convolutional layers often leads to prohibitive computational complexity and increased memory burden. Transformers, though effective in global dependency modeling, suffer from the quadratic complexity of their self-attention mechanism, which becomes a significant bottleneck for high-resolution SAR data. Furthermore, both architectures typically require extensive hyperparameter tuning, custom architectural adjustments, and protracted retraining cycles [59, 38], which can undermine their generalizability and robustness against SAR-specific challenges like speckle noise and significant resolution variations [21]. This analysis underscores a pressing need for novel, inherently lightweight architectural designs that possess stronger local feature extraction capabilities, superior parameter efficiency, and enhanced nonlinear expressive power beyond the conventional paradigms.
@@ II-B2 Application of KAN in Remote Sensing Field
Minjong Cheon [3] integrated KAN with various pre-trained CNN models to obtain KCN for remote sensing (RS) scene classification tasks. [52] proposed SpectralKAN, which can achieve high hyperspectral image change detection (HSIs-CD) accuracy, while requiring less parameters, FLOPs, GPU memory, training and testing time, and improving the efficiency of the HSIs-CD. [42] proposed a Wavelet-based Kolmogorov-Arnold Network Architecture (WaveKAN) that incorporates Wavelet functions into learnable activation functions for hyperspectral image classification. [39] integrated the KAN layer into the U-Net architecture (U-KAN), segmented farmland using Sentinel-2 and Sentinel-1 satellite images, and analyzed the performance and interpretability of these networks. Table II is a summary of the above methods.
@@ III-A The Design Motivation of Light-ResKAN
SAR images exhibit distinct characteristics compared to optical imagery, owing to its unique imaging mechanism and data properties [12]. Specifically, SAR images are severely afflicted by coherent speckle noise during the imaging process. Furthermore, due to the influence of side-looking imaging geometric distortions, their spatial resolution is typically lower, commonly ranging between 0.1 and 10 meters [36], [64]. This resolution is notably insufficient when contrasted with optical imagery acquired from the same platform.
@@ III-C1 Shared Activation Function Convolutions
As an orthogonal polynomial sequence, Gram Polynomials exhibit theoretical properties that align exceptionally well with the unique challenges of SAR imagery, positioning them as a highly promising activation function for SAR image recognition. Their principal advantage lies in their ability to adaptively integrate the processes of feature extraction and speckle noise suppression through adjustable polynomial degrees—effectively mitigating coherent speckle in homogeneous regions while enhancing critical features at edges and strong scatterers. Meanwhile, they adeptly handle the high dynamic range inherent in SAR data with their oscillatory characteristics. The orthogonality of their basis also contributes to numerical stability and generalization capability. More importantly, as demonstrated by the pioneering work of Liu et al. [27] in texture classification, polynomial like features can effectively capture the texture structure information of images. In addition, research on successful texture classification from random features [26] also indicates that methods based on function approximation can demonstrate significant strong potential in texture representation tasks. Consequently, from a theoretical perspective, Gram Polynomials offer a more powerful and intelligent nonlinear transformation compared to fixed functions like ReLU, providing a novel solution for SAR image recognition that combines mathematical rigor with enhanced practical adaptability.
@@ IV-B Experimental Setup
AlexNet [19] 2012 57.04 228.18 97.44 89.41 95.76 0.19 3.71 0.19 9.54 30.32 9.53
VGG16 [44] 2015 134.30 537.21 96.59 90.73 96.17 4.07 80.37 4.07 11.42 43.72 11.40
ResNet18 [13] 2016 11.18 44.73 96.71 83.54 95.34 0.49 9.51 0.49 10.31 32.41 10.29
ResNet50 [13] 2016 23.53 94.11 97.13 86.71 96.69 1.08 21.35 1.08 10.92 32.93 10.91
ConvNeXt [34] 2022 87.58 350.23 96.04 77.86 95.83 4.07 80.53 4.07 17.23 53.25 17.19
MACN [55] 2022 0.05 0.16 82.46 89.15 79.30 0.63 12.40 0.63 97.46 296.75 97.46
Cross-HL [41] 2024 0.52 2.06 98.19 78.06 97.03 0.01 0.22 0.01 11.57 35.64 11.56
ASANet [66] 2024 27.95 111.28 91.65 78.84 96.81 2.37 46.82 2.37 18.42 47.23 18.41
@@ IV-C Experimental Results
In this experiment, the performance of Light-ResKAN is comprehensively compared with a variety of representative models on the FUSAR-Ship, MSTAR, and SAR-ACD datasets. These baselines include: (1) Classic CNN architectures such as AlexNet [19], VGG16 [44], and the ResNet series (ResNet18, ResNet50) [13], which established the foundation for deep feature extraction. (2) ConvNeXt [34], a modernized pure CNN that integrates Transformer-like design principles. (3) Lightweight models specifically designed for efficiency, including MACN [55] which uses depthwise separable convolutions and Cross-HL [41] featuring cross-layer feature fusion. (4) ASANet [66], a multi-scale attention network specialized for SAR and RGB image recognition. (5) KAN-based competitors including SpectralKAN [52], WaveKAN [42], and the original ResKAGNet [6] which serves as the direct baseline for our proposed lightweight framework. Table VI shows the test accuracy, parameter quantity, model size, FLOPs per batch, and test time per batch of different models on three SAR dataset. The best results in each column (highest accuracy, lowest parameters, model size, FLOPs, and test time) are highlighted in bold.
@@
Test time. AlexNet yields the shortest test time due to its simple and highly optimized operator pipeline. Despite Light-ResKAN achieving superior accuracy with significantly reduced parameters and FLOPs, it currently exhibits higher inference latency compared to ResKAGNet and other lightweight baselines. This discrepancy highlights the disconnect between theoretical complexity metrics and actual runtime speed, a phenomenon comprehensively analyzed in ShuffleNet-V2 [37]. The observed latency primarily stems from two critical factors: the inherent computational characteristics of KANs and the specific engineering constraints necessitated by our implementation.
@@ V Discussions
AlexNet [19] 99.38 15.01 86.31
VGG16 [44] 99.34 320.83 131.50
ResNet18 [13] 99.05 38.01 90.47
ResNet50 [13] 99.17 86.58 101.47
ConvNeXt [34] 98.76 321.47 166.58
MACN [55] 92.53 49.48 956.23
Cross-HL [41] 99.43 0.84 89.92
ASANet [66] 95.36 186.81 145.93
@@ V-C1 Noise Characteristics of SAR Images
SAR systems utilize coherent radar technology, transmitting coherent pulses and performing coherent superposition of the echo signals upon reception [2]. When the radar beam illuminates the ground, a resolution cell contains multiple scatterers. The echoes from these scattering points are superimposed vectorially at the receiver. Due to the independent and random distribution of amplitudes and phases of the individual scatterers, the superimposed total echo intensity exhibits random fluctuations on top of the scattering coefficient, forming alternating bright and dark speckle noise. This superposition process directly leads to the noise being correlated with the local intensity of the signal, manifesting as multiplicative noise [49]. The specific formula is as follows:
@@
Where $I$ represents the observed value, $R$ represents the ideal radar scattering cross-section unaffected by noise, reflecting the backscattering characteristics of the ground object, and $F$ represents the coherent speckle noise component. To suppress noise, SAR often employs multi-look processing, which involves averaging multiple independent observations of the same area [9]. According to the Central Limit Theorem, the intensity noise after multi-look processing obeys a Gamma distribution, and the probability density function of its noise variable $F$ is:
@@
Where $\alpha$ is the shape parameter that determines the distribution shape, and $\beta$ is the scale parameter that controls the distribution’s scale. Additionally, studies have shown that the histogram of SAR image intensity data fits the Gamma distribution better than the Gaussian distribution, especially in low-resolution or complex scenes, where the gamma model can more accurately describe the clutter statistical properties [51]. Based on this, we add multiplicative noise following a Gamma distribution to three SAR datasets.